# Rotational mass of anisotropic neutron stars within Rastall gravity

### v2

@@ 1 Introduction
It is widely known that the mass of static neutron stars (NSs) can be obtained by solving Tolman-Oppenheimer-Volkoff (TOV) equation [1]. When we proceed to study the rotating NSs, we have to calculate the mass correction which arises in the formulation. Rotational mass of rotating relativistic stars within general relativity (GR) was proposed by Hartle & Thorne [2, 3]. In Hartle-Thorne (HT) formalism, rotation is handled as a perturbation for static configuration of the stars. As a direct consequence, the metric and the energy-momentum tensor (EMT) in HT formalism are expanded in multipole expansion, which is truncated at quadrupole order. HT formalism has been largely used by many authors (see Refs. [4, 5, 6, 7, 8, 9, 10, 11, 12, 13]), mostly concerning NSs [4, 5, 6, 7, 14, 15, 16]. Berti et al. [14] showed that HT approximation gives very good predictions for the corotating and counter-rotating radii $R_{\pm}$ of the innermost stable circular orbit (ISCO), with $R_{\pm}$ accurate to better than 1 per cent even for the fastest millisecond pulsars. It implies that HT formalism is relatively better than Manko formalism [17] and Cook-Shapiro-Teukolsky formalism [18], where both also deal with rotating NS.
@@ 2 Formalisms
In this section, we divide the formalism into three parts. In sub-section 2.1, we briefly review the Rastall gravity formalism. The anisotropic NSs’ rotational mass formalism within Hartle-Thorne-Rastall (HTR) framework is given in sub-section 2.2. The profile of anisotropic pressure is described in sub-section 2.3. It is worth noting that we use geometrized units $G=c=1$ in this paper.
@@ 2.1 Rastall gravity
From Eq. (4), we could take an insight that $\kappa\lambda=1/4$ mathematically gives us a singularity, so this value is forbidden. From Eq. (7) and (8), we can see that $\Lambda=-1/2$ and $\Lambda=1/4$ are also forbidden. We can verify that, as we expect, Rastall gravity recovers to GR when we set $\lambda=0$.
@@ 2.2 Rotational mass of anisotropic neutron stars
Here the analytical result of the rotational mass of NSs within HTR formalism which is generalized to anisotropic pressure conditions is presented. First of all, we recall the metric of rotating star, which writes [4]
$ds^{2}$ $=$ $-e^{2\nu}dt^{2}+e^{2\lambda}dr^{2}+r^{2}e^{2\psi}(d\phi-\omega\>dt)^{2}$ (9)
$+r^{2}e^{2\mu}d\theta^{2},$
where $\omega$ denotes the angular velocity of the local inertial frame, which is proportional to the star’s angular velocity $\Omega$ relative to a distant observer. Both $\omega$ and $\Omega$ satisfy $\omega=\Omega-\bar{\omega}$, where $\bar{\omega}$ is the angular velocity of the star relative to the local inertial frame.
In standard HT procedure, the metric functions are expanded as follows [2, 3, 4]
$e^{2\nu}$ $=$ $e^{2\varphi}[1+2(h_{0}+h_{2}P_{2}(\cos\theta))],$ (10)
$e^{2\lambda}$ $=$ $\left[1+\frac{2}{r}\left(m_{0}+m_{2}P_{2}(\cos\theta)\right)\left(1-\frac{2m(r)}{r}\right)^{-1}\right]$ (11)
×$\times\left(1-\frac{2m(r)}{r}\right)^{-1}$
$e^{2\psi}$ $=$ $r^{2}\sin^{2}\theta[1+2(v_{2}-h_{2})P_{2}(\cos\theta)],$ (12)
$e^{2\mu}$ $=$ $r^{2}[1+2(v_{2}-h_{2})P_{2}(\cos\theta)].$ (13)
Here $h_{0}$, $h_{2}$, $m_{0}$, $m_{2}$, and $v_{2}$ are perturbation functions; $P_{2}(\cos\theta)$ is the second order of Legendre polynomial equation; $e^{2\varphi}$ is a function which is in GR constrained by
$\left(\frac{d\varphi}{dr}\right)_{GR}=\frac{m+4\pi r^{3}p}{r(r-2m)},$ (14)
where $p$ is the pressure and $m$ is mass fuction which is in GR constrained by
$\left(\frac{dm}{dr}\right)_{GR}=4\pi r^{2}\varepsilon.$ (15)
Here $\varepsilon$ denotes the energy density.
Based on Refs. [15, 19, 71], the EMT of anisotropic NSs could be written as
$T^{\mu}_{\nu}=(\varepsilon+q)u^{\mu}u_{\nu}+q\delta^{\mu}_{\nu}+\sigma\chi^{\mu}\chi_{\nu},$ (16)
where $u_{\nu}$ is the fluid 4-velocity, $\chi_{\nu}$ is a unit radial vector that satisfies $u^{\mu}\chi_{\nu}=0$, $q$ is the tangential pressure which satisfies $\sigma=p-q$, where $p$ is the radial pressure. It is important to note that it is valid to assume that $T^{r}_{\theta}=0$. Such assumption has been used in some studies regarding non-rotating anisotropic NS (please see Ref. [73, 74]) and rotating anisotropic NSs (please see Refs. [15, 16, 75]). Considering Eq. (6), we can write
$\tilde{T}^{\mu}_{\nu}=(\tilde{\varepsilon}+\tilde{q})u^{\mu}u_{\nu}+\tilde{q}\delta^{\mu}_{\nu}+\sigma\chi^{\mu}\chi_{\nu},$ (17)
where $\tilde{\varepsilon}=\varepsilon+\left(\frac{\kappa\lambda}{4\kappa\lambda-1}\right)T$, $\tilde{p}=p-\left(\frac{\kappa\lambda}{4\kappa\lambda-1}\right)T$, and $\tilde{q}=q-\left(\frac{\kappa\lambda}{4\kappa\lambda-1}\right)T$.
The Rastall version of Eq. (14) reads
@@
while the Rastall version of Eq. (15) reads
@@
When we impose Rastall gravity on anisotropic TOV equation, we can obtain the Rastall version of the equation as follows
@@
We have to note that TOV equation is based on static condition. For a rotating system, we have to consider the angular velocity relative to the local inertial frame, which can be obtained by numerical calculation on the following equation
$\frac{1}{r^{4}}\frac{d}{dr}\left(r^{4}j\frac{d\bar{\omega}}{dr}\right)+\frac{4}{r}\frac{dj}{dr}\left(1-\frac{\sigma}{\varepsilon+p}\right)=0,$ (21)
where
$j$ $=$ $e^{-\varphi}\left(1-\frac{2m}{r}\right)^{1/2},$ (22)
and boundary conditions at $r=0$ are $\bar{\omega}=\bar{\omega}_{c}$ and $\frac{d\bar{\omega}}{dr}=0$; where $\bar{\omega}_{c}$ can be chosen arbitrarily. Furthermore, it is important to note that Eq. (21) is not modified due to Rastall gravity.
The corrections of Einstein tensor within HTR formalism is given by
$\Delta G^{\mu}_{\nu}=\delta G^{\mu}_{\nu}-\kappa\xi\frac{d}{dr}\tilde{T}^{\mu}_{\nu},$ (23)
where $\delta G^{\mu}_{\nu}$ denotes the expansion terms of $G^{\mu}_{\nu}$. In the original HT formalism, the expansion of $\xi$ is limited to $l=2$, and is given by
$\xi=\xi_{0}(r)+\xi_{2}(r)P_{2}(\cos\theta),$ (24)
where
$\xi_{0}=-\mathcal{P}_{0}\left(\varepsilon+p\right)\left(1-\frac{\sigma}{\varepsilon+p}\right)\left(\frac{dp}{dr}-\frac{d\sigma}{dr}\right)^{-1},$ (25)
$\xi_{2}=-\mathcal{P}_{2}\left(\varepsilon+p\right)\left(1-\frac{\sigma}{\varepsilon+p}\right)\left(\frac{dp}{dr}-\frac{d\sigma}{dr}\right)^{-1}.$ (26)
Here $\mathcal{P}_{0}$ and $\mathcal{P}_{2}$ are expanded terms of pressure perturbation factors $\mathcal{P}$ which are given by
$\mathcal{P}=\mathcal{P}_{0}+\mathcal{P}_{2}\>P_{2}(\cos\theta).$ (27)
The Rastall field equation of the tensors’ corrections reads,
$\Delta G^{\mu}_{\nu}=\kappa\Delta\tilde{T}^{\mu}_{\nu}=\kappa\left[\Delta T^{\mu}_{\nu}-\delta^{\mu}_{\nu}\left(\frac{\kappa\lambda}{4\kappa\lambda-1}\Delta T\right)\right].$ (28)
The corrections of EMT $\Delta T^{\mu}_{\nu}$ are given by
$(\Delta T^{t}_{t})_{l=0}=-\frac{2}{3}\left(\varepsilon+p\right)e^{-2\varphi}r^{2}\bar{\omega}\left(1-\frac{\sigma}{\varepsilon+p}\right),$ (29)
$(\Delta T^{r}_{r})_{l=0}=(\Delta T^{\theta}_{\theta})_{l=0}=(\Delta T^{\phi}_{\phi})_{l=0}=0.$ (30)
It is clear that according to Eq. (28), we need to substitute the corrections of energy-momentum scalar $\Delta T$ in the calculation. Since the rotational mass only involves the $l=0$ order, we can just focus on this order. The corrections of energy-momentum scalar is given by
$\left(\Delta T\right)_{l=0}$ $=$ $-\frac{2}{3}\left(\varepsilon+p\right)e^{-2\varphi}r^{2}\bar{\omega}\left(1-\frac{\sigma}{\varepsilon+p}\right).$ (31)
@@
$\frac{dm_{0}}{dr}$ $=$ $\frac{\kappa}{2}r^{2}\frac{d\tilde{\varepsilon}}{dp}\left(1-\frac{d\sigma}{dp}\right)^{-1}\left(\varepsilon+p\right)$ (32)
×$\times\left(1-\frac{\sigma}{\varepsilon+p}\right)\mathcal{P}_{0}+\frac{1}{12}j^{2}r^{4}\left(\frac{d\bar{\omega}}{dr}\right)^{2}$
$-\frac{1}{3}r^{3}\left(\frac{dj^{2}}{dr}\right)\bar{\omega}^{2}\left(1-\frac{\sigma}{\varepsilon+p}\right)$
@@
while from the ($rr$)-component of the $l=0$ order, one can obtain
$\frac{d\mathcal{P}_{0}}{dr}$ $=$ $-\frac{m_{0}\left(1+\kappa r^{2}\tilde{p}\right)}{\left(r-2m\right)^{2}}-\frac{\kappa r^{2}\left(\varepsilon+p\right)}{2\left(r-2m\right)}$ (33)
×$\times\left(1-\frac{\sigma}{\varepsilon+p}\right)\left(1-\frac{d\sigma}{dp}\right)^{-1}\beta\mathcal{P}_{0}$
$+\frac{1}{12}\frac{r^{4}j^{2}}{r-2m}\left(\frac{d\bar{\omega}}{dr}\right)^{2}$
@@
where
$\beta$ $=$ $1+\frac{\kappa\lambda}{4\kappa\lambda-1}\left(1-\frac{d\varepsilon}{dp}\right)+\frac{2\kappa\lambda}{4\kappa\lambda-1},$ (34)
$\gamma$ $=$ $\frac{r^{3}}{r-2m}\frac{\kappa\lambda}{4\kappa\lambda-1}\left(\frac{dj^{2}}{dr}\right)\bar{\omega}^{2}\left(1-\frac{\sigma}{\varepsilon+p}\right).$ (35)
Remember that both $m_{0}$ and $\mathcal{P}_{0}$ vanish at the origin. Moreover, as we expect, if the formalism recovers to GR (i.e. $\lambda=0$), then $\beta=1$ and $\gamma=0$, and Eq. (33) recovers to Eq. (62) in Ref. [16]. It has to be highlighted that there is a typo in Eq. (62) of Ref. [16]. The second term of the right-handed side should include the factor $\left(1-\frac{d\sigma}{dp}\right)^{-1}$. For the differential equations consisting $m_{0}$ and $\mathcal{P}_{0}$, the boundary conditions are $m_{0}(r=0)=0$ and $\mathcal{P}_{0}(r=0)=0$.
@@
From Eq. (36), the moment of inertia $I$ of NSs can be calculated
$I=\frac{J}{\Omega}.$ (37)
Mass correction of the stars is given by
$\delta M=m_{0}(r=R)+\frac{J^{2}}{R^{3}},$ (38)
where $R$ is radius of the stars.
Finally, the rotational mass reads

### v3

@@ 1 Introduction
It is widely known that the mass of static neutron stars (NSs) can be obtained by solving Tolman-Oppenheimer-Volkoff (TOV) equation [1]. When we proceed to study the rotating NSs, we have to calculate the mass correction which arises in the formulation. Rotational mass of rotating relativistic stars within general relativity (GR) was proposed by Hartle & Thorne [2, 3]. In Hartle-Thorne (HT) formalism, rotation is handled as a perturbation for static configuration of the stars. As a direct consequence, the metric and the energy-momentum tensor (EMT) in HT formalism are expanded in multipole expansion, which is truncated at quadrupole order. HT formalism has been largely used by many authors (see Refs. [4, 5, 6, 7, 8, 9, 10, 11, 12, 13]), mostly concerning NSs [4, 5, 6, 7, 14, 15, 16].
Berti et al. [14] conducted a comparative analysis of the HT approximation, the Cook–Shapiro–Teukolsky (CST) solutions [17], and the Manko–Miel-ke–Sanabria-Gómez (MMSG) formalism [18]. Note that the CST framework accommodates both slowly and rapidly rotating compact stars. In the comparison between the HT approximation and the CST models, configurations were selected such that the gravitational mass and angular momentum were matched. Their results demonstrate that the HT formalism provides remarkably accurate predictions for the corotating and counterrotating radii, $R_{\pm}$, of the innermost stable circular orbit, with deviations less than 1 % even for the most rapidly rotating millisecond pulsars. When the HT approximation is compared to the MMSG formalism—specifically in the exterior vacuum spacetime—it also demonstrates better overall performance and higher accuracy.
@@ 2 Formalisms
In this section, we divide the formalism four parts. In sub-section 2.1, we briefly review the Rastall gravity formalism. The nuclear model for the EOS is shortly discussed in sub-section 2.2. The anisotropic NSs’ rotational mass formalism within Hartle-Thorne-Rastall framework is given in sub-section 2.3. The profile of anisotropic pressure is described in sub-section 2.4. It is worth noting that we use geometrized units $G=c=1$ in this paper.
@@ 2.1 Rastall gravity
Ref. [73] provides an additional important informations regarding allowed values of the Rastall parameter. The $\kappa$ has to satisfy
${\kappa=\frac{4\kappa\lambda-1}{6\kappa\lambda-1}8\pi,}$ (9)
and the $\lambda$ has to satisfy
${\lambda=\frac{\kappa\lambda(6\kappa\lambda-1)}{(4\kappa\lambda-1)8\pi}.}$ (10)
From Eq. (4), we could take an insight that $\kappa\lambda=1/4$ mathematically gives us a singularity, so this value is forbidden. From Eq. (7) and (8), we can see that $\Lambda=-1/2$ and $\Lambda=1/4$ are also forbidden. Moreover, Eq. (9) conveys the same information as Eq. (4), while Eq. (10) reveals that the condition $\kappa\lambda=1/6$ is also not allowed. We can verify that, as we expect, Rastall gravity recovers to GR when we set $\lambda=0$.
2.2 Nuclear model of the EOS
The Lagrangian of the nuclear model used in this work is given by $\mathcal{L}=\mathcal{L}_{B}+\mathcal{L}_{M}+\mathcal{L}_{\text{int}}$ [68].
Here $\mathcal{L}_{B}$ denotes the baryon term which reads
${\mathcal{L}_{B}=\sum_{B}\bar{\psi}_{B}(i\gamma_{\mu}\partial^{\mu}-M_{B})\psi_{B},}$ (11)
where $\psi_{B}$ is the baryon field and $M_{B}$ is the baryon mass in vacuum. Note that the sum $B$ runs over the octet baryons $p$, $n$, $\Lambda$, $\Sigma^{+0-}$, and $\Xi^{0-}$. Here $M_{B}$ denotes baryon mass.
The Lagrangian of meson term $\mathcal{L}_{M}$ reads
${\mathcal{L}_{M}}$ $=$ ${\frac{1}{2}\left(\partial_{\mu}\sigma\,\partial^{\mu}\sigma-m_{\sigma}^{2}\sigma^{2}\right)}+{\frac{1}{2}m_{\omega}^{2}\omega_{\mu}\omega^{\mu}}-{\frac{1}{4}W_{\mu\nu}W^{\mu\nu}}$ (12)
${+\frac{1}{2}m_{\rho}^{2}\,\vec{\rho}_{\mu}\cdot\vec{\rho}^{\mu}}-{\frac{1}{4}\vec{R}_{\mu\nu}\cdot\vec{R}^{\mu\nu}}$
${+\frac{1}{2}\left(\partial_{\mu}\vec{\pi}\cdot\partial^{\mu}\vec{\pi}-m_{\pi}^{2}\vec{\pi}^{2}\right)}.$
where ${W_{\mu\nu}}={\partial_{\mu}\omega_{\nu}-\partial_{\nu}\omega_{\mu}}$ and ${\vec{R}_{\mu\nu}}={\partial_{\mu}\vec{\rho}_{\nu}-\partial_{\nu}\vec{\rho}_{\mu}.}$ Here $\sigma$ and $\omega$ are isoscalar mesons, and $\vec{\pi}$ and $\vec{\rho}$ are isovector meson.
The lagrangian of interaction $\mathcal{L}_{\text{int}}$ writes
${\mathcal{L}_{\text{int}}}$ $=$ ${\sum_{B}\bar{\psi}_{B}\Bigg{[}g_{\sigma B}(\sigma)\sigma-g_{\omega B}\gamma_{\mu}\omega^{\mu}+\frac{f_{\omega B}}{2\mathcal{M}}\sigma_{\mu\nu}\partial^{\nu}\omega^{\mu}}$ (13)
${-g_{\rho B}\gamma_{\mu}\,\vec{\rho}^{\mu}\cdot\vec{I}_{B}+\frac{f_{\rho B}}{2\mathcal{M}}\sigma_{\mu\nu}\partial^{\nu}\vec{\rho}^{\mu}\cdot\vec{I}_{B}}$
${-\frac{f_{\pi B}}{m_{\pi}}\gamma_{5}\gamma_{\mu}\,\partial^{\mu}\vec{\pi}\cdot\vec{I}_{B}\Bigg{]}\psi_{B}.}$
Here the common scale mass $\mathcal{M}$ is taken to be free nucleon mass, the commutation operator for the $\gamma$ matrix writes $\sigma_{\mu\nu}=i\frac{[\gamma_{\mu},\gamma_{\nu}]}{2}$, $\vec{I}_{B}$ denotes the isospin matrix for baryon $B$, which is set to $\vec{I}_{B}=0$ for iso-singlet baryons. Note that $f_{\omega B}$ and $f_{\rho B}$ are isoscalar-tensor and isovector-tensor coupling constants, respectively. the terms involving $g_{\sigma B}(\sigma)$, $g_{\omega B}$, $g_{\rho B}$, and $f_{\pi B}$ correspond to $\sigma-$, $\omega-$, $\rho-$, and $\pi-B$ couplings, respectively. Also note that $g_{\omega B}$, $g_{\rho B}$, $f_{\omega B}$, $f_{\rho B}$, and $f_{\pi B}$ are constants. In this formulation, the meson field values are replaced by the constant mean-field values, i.e. $\bar{\sigma}$, $\bar{\omega}$, and $\bar{\rho}$ (the $\rho^{0}$ field).
$B$ $N$ $\Lambda$ $\Sigma$ $\Xi$
$g_{\sigma B}/\sqrt{4\pi}$ 1.94 2.15 1.67 1.50
$g_{\omega B}/\sqrt{4\pi}$ 2.39 2.82 2.82 2.09
$f_{\omega B}/\sqrt{4\pi}$ $-0.545$ $-3.39$ $-0.261$ $-4.40$
$g_{\rho B}/\sqrt{4\pi}$ 0.596 0 1.19 0.596
$f_{\rho B}/\sqrt{4\pi}$ 3.39 0 2.94 $-0.446$
$f_{\pi B}/\sqrt{4\pi}$ 0.268 0 $-0.0772$
Table 1: Coupling constants used in this model.
The coupling constants used in this model is shown by Table 1, while the prediction of the model at saturation is presented in Table 2.
$w_{0}$ $n_{0}$ $S_{0}$ $K_{0}$ $L$
(MeV) (fm-3) (MeV) (MeV) (MeV)
${-16.1}$ 0.1550.155{0.155}0.155 33.633.6{33.6}33.6 274274{274}274 77.177.1{77.1}77.1
Table 2: Predictions of this model at saturation density, including saturation energy $w_{0}$, saturation density $n_{0}$ itself, symmetry energy $S_{0}$, incompressibility $K_{0}$, and slope of the symmetry energy $L$.
In this model [68], at the inner crust, a fraction of neutrons become unbound and ”drip” out of the nuclei; while at the outer crust, neutrons remain bound within the nuclei. In the core, there exist hyperons, where the potential depth of the hyperons reads
$U_{Y}^{(N)}$ $=$ ${-g_{\sigma Y}(\bar{\sigma}^{(N)})\bar{\sigma}^{(N)}+g_{\omega Y}\bar{\omega}^{(N)}}$ (14)
${+\frac{1}{2M_{Y}}\left[-g_{\sigma Y}(\bar{\sigma}^{(N)})\bar{\sigma}^{(N)}+g_{\omega Y}\bar{\omega}^{(N)}\right]^{2}.}$
The hyperon species $\Lambda$, $\Sigma$, and $\Xi$ are collectively represented by $Y$. The quantities $\bar{\sigma}^{(N)}$ and $\bar{\omega}^{(N)}$ correspond to the mean-field values of the $\sigma$ and $\omega$ meson fields in symmetric nuclear matter, respectively. The details of this model—including the way to obtain the energy density—can be referred to Ref. [68]. After one obtain the energy density $\varepsilon$, the pressure $p$ can be calculated as $p=n_{B}^{2}\frac{\partial}{\partial n_{B}}\left(\frac{\varepsilon}{n_{B}}\right),$ where $n_{B}$ denotes the total baryon number density. It is evident that the EOS model employed in this work incorporates physically realistic components
@@ 2.3 Rotational mass of anisotropic neutron stars
The metric of rotating NSs with HT approximation is written as [4]
$ds^{2}$ $=$ $-e^{2\varphi}[1+2(h_{0}+h_{2}P_{2}(\cos\theta))]dt^{2}$ (15)
$+\left[1+\frac{2}{r}\left(m_{0}+m_{2}P_{2}(\cos\theta)\right)\left(1-\frac{2m(r)}{r}\right)^{-1}\right]\left(1-\frac{2m(r)}{r}\right)^{-1}dr^{2}$
$+\>r^{2}\sin^{2}\theta[1+2(v_{2}-h_{2})P_{2}(\cos\theta)](d\phi-\omega\>dt)^{2}$
$+r^{2}[1+2(v_{2}-h_{2})P_{2}(\cos\theta)]d\theta^{2},$
Here, $\omega$ denotes the angular velocity of the local inertial frame, which is proportional to the star’s angular velocity $\Omega$ relative to a distant observer. Both $\omega$ and $\Omega$ satisfy $\omega=\Omega-\bar{\omega}$, where $\bar{\omega}$ is the angular velocity of the star relative to the local inertial frame. Note that $h_{0}$, $h_{2}$, $m_{0}$, $m_{2}$, and $v_{2}$ are perturbation functions; $P_{2}(\cos\theta)$ is the second order of Legendre polynomial equation; $e^{2\varphi}$ is a metric function.
The anisotropic form of the effective EMT writes
$\tilde{T}^{\mu}_{\nu}=(\tilde{\varepsilon}+\tilde{q})u^{\mu}u_{\nu}+\tilde{q}\delta^{\mu}_{\nu}+\sigma\chi^{\mu}\chi_{\nu}.$ (16)
Here $u_{\nu}$ is the fluid 4-velocity, $\chi_{\nu}$ is a unit radial vector that satisfies $u^{\mu}\chi_{\nu}=0$, $q$ is the tangential pressure which satisfies $\sigma=p-q$, where $p$ is the radial pressure, $\tilde{\varepsilon}=\varepsilon+\left(\frac{\kappa\lambda}{4\kappa\lambda-1}\right)T$, $\tilde{p}=p-\left(\frac{\kappa\lambda}{4\kappa\lambda-1}\right)T$, and $\tilde{q}=q-\left(\frac{\kappa\lambda}{4\kappa\lambda-1}\right)T$. It is important to note that it is valid to assume that $T^{r}_{\theta}=0$ [15, 16, 74, 75, 76].
The function $e^{2\varphi}$ is constrained by
@@
while the mass $m$ is constrained by
@@
The anisotropic TOV equation within Rastall gravity reads
@@
For rotating relativistic stars, the $\bar{\omega}(r)$, with an initial arbitrary central value $\bar{\omega}(r=0)=\omega_{c}$, can be determined by solving Eq. (72) in Ref. [16].
@@
$\frac{dm_{0}}{dr}$ $=$ $\frac{\kappa}{2}r^{2}\frac{d\tilde{\varepsilon}}{dp}\left(1-\frac{d\sigma}{dp}\right)^{-1}\left(\varepsilon+p\right)\left(1-\frac{\sigma}{\varepsilon+p}\right)\mathcal{P}_{0}$ (20)
$+\frac{1}{12}j^{2}r^{4}\left(\frac{d\bar{\omega}}{dr}\right)^{2}-\frac{1}{3}r^{3}\left(\frac{dj^{2}}{dr}\right)\bar{\omega}^{2}\left(1-\frac{\sigma}{\varepsilon+p}\right)$
@@
where $j=e^{-\varphi}\left(1-\frac{2m}{r}\right)^{1/2}$. From the ($rr$)-component of the $l=0$ order, one can obtain
$\frac{d\mathcal{P}_{0}}{dr}$ $=$ $-\frac{m_{0}\left(1+\kappa r^{2}\tilde{p}\right)}{\left(r-2m\right)^{2}}-\frac{\kappa r^{2}\left(\varepsilon+p\right)}{2\left(r-2m\right)}\left(1-\frac{\sigma}{\varepsilon+p}\right)$ (21)
×$\times\left(1-\frac{d\sigma}{dp}\right)^{-1}\beta\>\mathcal{P}_{0}+\frac{1}{12}\frac{r^{4}j^{2}}{r-2m}\left(\frac{d\bar{\omega}}{dr}\right)^{2}$
@@
where $\beta=1+\frac{\kappa\lambda}{4\kappa\lambda-1}\left(1-\frac{d\varepsilon}{dp}\right)+\frac{2\kappa\lambda}{4\kappa\lambda-1}$ and $\gamma=\frac{r^{3}}{r-2m}\frac{\kappa\lambda}{4\kappa\lambda-1}\left(\frac{dj^{2}}{dr}\right)\bar{\omega}^{2}\left(1-\frac{\sigma}{\varepsilon+p}\right).$ Here $\mathcal{P}_{0}$ is the pressure perturbation factor. Remember that $m_{0}(r=0)=\mathcal{P}_{0}(r=0)=0$.
@@
From Eq. (22), the moment of inertia $I$ of NSs can be calculated, i.e. $I=\frac{J}{\Omega}$.
Mass correction of the stars is given by $\delta M=m_{0}(r=R)+\frac{J^{2}}{R^{3}}$, where $R$ is radius of the stars. Finally, the rotational mass reads